You have determined that you will need $1,500,000 to fund your retirement. If you plan to retire in 30 years, and you believe you can earn a return of 9%, how much must you invest each month in order to achieve your goal?
Answer #1

Interest per month is (9%/12) = 0.75%.

The formula is given below:

Future\, value=Monthly\, payment\times \left ( \frac{\left ( 1+i \right )^{n}-1}{i} \right )

1,500,000=Monthly\, payment\times \left ( \frac{\left ( 1+0.0075 \right )^{12\times 30}-1}{0.0075} \right )

1,500,000=Monthly\, payment\times \left ( \frac{14.73057612-1}{0.0075} \right )

1,500,000=Monthly\, payment\times 1,830.743483\right )

819.34=Monthly\, payment